# Multi-$\text{P}^{\text{2}}$: A Multi-perspective Benchmark on Privacy Assessment for Large Vision-Language Models

### v2

@@ Appendix A Discussion
A.1 Concerns about Privacy Image Recognition
A.2 Limitation
Multi-$\text{P}^{\text{2}}$ focuses on evaluating the privacy preservation capabilities of Large Vision-Language Models (LVLMs) with respect to image inputs, lacking corresponding tests for textual inputs. Although existing privacy benchmarks for large language models (LLMs) have assessed privacy concerning textual inputs, the limitations in privacy types, and the uncertainties introduced in model outputs due to the incorporation of image inputs in LVLMs, warrant further investigation into the privacy risks of textual information by LVLMs. We plan to address this gap in future work.
Furthermore, we do not evaluate the model’s privacy preservation capabilities under privacy attack scenarios. Recent studies have demonstrated that jailbreak attacks can bypass model safety mechanisms, leading to unsafe outputs. Since jailbreak attacks do not require prior knowledge of the model’s training data and can easily be implemented through prompt engineering to extract private information, the need for privacy evaluation under jailbreak attack scenarios is both crucial and urgent. We plan to incorporate diverse jailbreak attack scenarios in subsequent iterations of our benchmark to provide a more comprehensive evaluation of models’ privacy preservation capabilities.
Finally, privacy exhibits temporal sensitivity, meaning that over time, our privacy evaluation samples may become inadequate to assess future large vision-language models. This time-dependent nature of privacy implies that evaluation samples require continuous updates as models progress. To address this, we plan to develop a dynamic and scalable privacy assessment framework capable of adapting to the advancements of LVLMs over time.
A.3 Construction of Sensitive Questions
@@
A.4 Social and Ethical Concerns
The privacy categories encompassed by Multi-$\text{P}^{\text{2}}$ are formulated by a team of researchers with expertise in computer science and psychology, and offer valuable insights for evaluating the privacy preservation capabilities of Large Vision-Language Models (LVLMs). Notably, trade secret and state secret are considered highly sensitive types of confidential information. To mitigate the risk of attackers leveraging our benchmark to extract such sensitive data, we rigorously filtered the collected images, retaining only publicly available, low-sensitivity image data related to trade and nation. We believe evaluating privacy risks LVLMs on these two privacy types is essential, as a model’s ability to protect highly sensitive data reflects its potential societal risks and demonstrates responsible consideration for its deployment and application. Multi-$\text{P}^{\text{2}}$ may empower model developers to gain a more granular understanding of their models’ privacy vulnerabilities and implement appropriate safety mechanisms to mitigate these risks, ultimately contributing to the development of privacy-enhanced LVLMs.
A.5 Copyright Attribution for Images
@@ Appendix B Information about Dataset and Models
B.1 Dataset Statistics
In total, Multi-$\text{P}^{\text{2}}$ comprises 31,962 samples. Statistics for each task in Multi-$\text{P}^{\text{2}}$ are presented in Table 5.
@@ B.1 Dataset Statistics
Table 5: Statistics of Multi-$\text{P}^{\text{2}}$.
B.2 Model Hub
Configurations and specifications for some open-source models evaluated in our experiments are listed in Table 6.
model VE LLM blip2-opt-3b [22] ViT-g/14(EVA-CLIP) OPT-3B blip2-opt-7b [22] ViT-g/14(EVA-CLIP) OPT-7B blip2_flan-t5-xl [22] ViT-g/14(EVA-CLIP) FlanT5-XL instructblip_flan-t5-xl [13] ViT-g/14(EVA-CLIP) FlanT5-XL instructblip_flan-t5-xxl [13] ViT-g/14(EVA-CLIP) FlanT5-XXL instructblip_vicuna-13b [13] ViT-g/14(EVA-CLIP) Vicuna-13B internlm-xcomposer-vl-7b [51] ViT-g/14(EVA-CLIP) internlm-xcomposer-7b llava_1.5-13b [25] ViT-L/14-336(CLIP) Vicuna-13B llava_1.5-7b [25] ViT-L/14-336(CLIP) Vicuna-7B minigpt4_llama2 [54] BLIP2-VE(EVA-CLIP) LLaMA2 minigpt4_vicuna-13b [54] BLIP2-VE(EVA-CLIP) Vicuna-13B minigpt_v2 [11] ViT-g/14(EVA-CLIP) LLaMA2 mplug-owl2 VIT-L(CLIP) LLaMA otter [20] ViT-L/14(CLIP) LLaMA qwen-vl-chat [6] ViT-G(CLIP) Qwen-7b shikra-7b [12] ViT-L/14(CLIP) Vicuna-7B
Table 6: Configurations of some open-source LVLMs evaluated in our experiments, “VE” stands for visual encoder and “LLM” stands for language model.
@@ Appendix C Details of dataset generation
C.1 Detailed Privacy Categories
@@
C.2 Questions of Privacy Question Detection
@@
C.3 Questions of Privacy InfoFlow Assessment
@@
C.4 Templates of Safety Enhancement
@@
Appendix D Additional Experiment Results
D.1 Details of Privacy Question Detection
Model $ACC_{Perception}$ $ACC_{Reasoning}$ $ACC_{Memory}$ $ACC_{Insensitive}$ blip2-opt-3b 0.7329 0.6331 0.5798 0.2507 blip2-opt-7b 0.2993 0.2625 0.2333 0.2039 blip2-flan-t5-xl 0.0054 0.0000 0.0158 0.9254 glm-4v-9b 0.0663 0.0127 0.0548 0.9203 instructblip_flan-t5-xl 0.0143 0.0000 0.0205 0.9254 instructblip_flan-t5-xxl 0.0493 0.0142 0.0585 0.9254 instructblip_vicuna-13b 0.0556 0.1767 0.0843 0.8847 instructblip_vicuna-7b 0.4642 0.5831 0.4123 0.5304 internlm-xcomposer-vl-7b 0.0376 0.0089 0.0595 0.9243 llava_1.5-13b 0.5824 0.3989 0.4212 0.6488 llava_1.5-7b 0.6801 0.5652 0.6319 0.4309 minicpm_llama2-v2.5 0.3029 0.2132 0.2791 0.6246 minigpt4_llama_2 0.1344 0.1074 0.1253 0.5840 minigpt4_vicuna-13b 0.0779 0.0753 0.0737 0.2390 minigpt_v2 0.5761 0.4817 0.4787 0.5453 mplug-owl2 0.2384 0.1059 0.2275 0.8498 otter 0.8118 0.5645 0.5350 0.4989 phi-3-vision 0.4597 0.2960 0.3239 0.8535 qwen-vl-chat 0.3334 0.2491 0.2823 0.7734 shikra-7b 0.1353 0.0999 0.1201 0.4287 yi-vl 0.0099 0.0000 0.0263 0.9254 GPT-4o 0.6989 0.4932 0.6040 0.7971 Gemini-1.5-pro 0.5418 0.4209 0.5675 0.8584
Table 7: Detailed results on Privacy Question Detection.
As shown in Table 7, for closed-source models like GPT-4o [33] and Gemini-1.5-pro [36], accurately identifying questions related to privacy reasoning presents a significant challenge, where models exhibit poorer performance in Reasoning Leakage compared to other types of privacy leakage. For open-source models, the results for Privacy Question Detection vary considerably across different tasks. Some models show a large disparity in accuracy between sensitive and insensitive questions. For example, blip2-opt-3b [22] tends to select “Yes” in judgment-type questions. Since the “Yes” option indicates that the model perceives the question as involving sensitive information, this tendency results in high accuracy for detecting sensitive questions but poor performance on insensitive ones. Conversely, some models, such as blip2-opt-7b [22], perform poorly on both sensitive and non-sensitive questions. This is often due to the model’s inability to effectively handle judgment questions, frequently failing to provide the necessary options (Yes or No) in its responses.
D.2 Question Detection of different privacy types
Model Personal Privacy Trade Secret State Secret blip2-opt-3b 0.4401 0.4668 0.4378 blip2-opt-7b 0.2119 0.3464 0.2153 blip2-flan-t5-xl 0.4387 0.5043 0.5084 glm-4v-9b 0.4678 0.5058 0.5020 instructblip_flan-t5-xl 0.4387 0.5154 0.5110 instructblip_flan-t5-xxl 0.4575 0.5222 0.5215 instructblip_vicuna-13b 0.4650 0.5237 0.5456 instructblip_vicuna-7b 0.4696 0.5488 0.5490 internlm-xcomposer-vl-7b 0.4489 0.5265 0.5284 llava_1.5-13b 0.5421 0.5616 0.5700 llava_1.5-7b 0.5068 0.5369 0.5664 minicpm_llama2-v2.5 0.4490 0.4874 0.4119 minigpt4_llama_2 0.3244 0.4207 0.3802 minigpt4_vicuna-13b 0.1682 0.1356 0.1454 minigpt_v2 0.5071 0.5622 0.5466 mplug-owl2 0.5009 0.5507 0.5501 otter 0.5353 0.6009 0.5826 phi-3-vision 0.6039 0.6136 0.5898 qwen-vl-chat 0.5108 0.5767 0.5458 shikra-7b 0.2710 0.2864 0.2727 yi-vl 0.4451 0.5043 0.5051 GPT-4o 0.6839 0.7128 0.7115 Gemini-1.5-pro 0.6683 0.6969 0.7236
Table 8: Detailed results on Privacy Question Detection. ACC is reported for evaluation. Samples of each privacy type consist of sensitive questions and insensitive questions.
Overall, most models exhibit relatively consistent accuracy across three privacy types (see Table 8). GPT-4o [33] outperforms all other models, achieving the highest accuracy in personal privacy and trade secret, indicating its superior capability in identifying privacy sensitivity of input queries. Gemini-1.5-pro [36] also demonstrates strong performance, slightly trailing GPT-4o [33], particularly in State Secret, where its accuracy reaches 0.7236. Current models exhibit similar performance in recognizing the privacy sensitivity of questions across three privacy types, whereas their performance varies significantly in Privacy Leakage among these privacy types. This disparity further suggests that current privacy preservation mechanisms of LVLMs do not enhance their security at a conceptual or awareness level, highlighting potential limitations in their ability to safeguard privacy effectively.
D.3 Question Detection in Safety-enhanced Setting
Model Sensitive Questions Insensitive Questions w/o w/ w/o w/ blip2-opt-3b 0.6486 0.5993 0.2507 0.3480 blip2-opt-7b 0.2650 0.3943 0.2039 0.4444 blip2_flan-t5-xl 0.0071 0.0468 0.9254 0.9254 glm-4v-9b 0.0446 0.0678 0.9203 0.9240 instructblip_flan-t5-xl 0.0116 0.0540 0.9254 0.9251 instructblip_flan-t5-xxl 0.0406 0.0851 0.9254 0.9254 instructblip_vicuna-13b 0.1055 0.0219 0.8849 0.0249 instructblip_vicuna-7b 0.4865 0.4224 0.5303 0.1915 internlm-xcomposer-vl-7b 0.0354 0.0416 0.9243 0.9232 llava_1.5-13b 0.4676 0.2330 0.6488 0.8439 1lava_1.5-7b 0.6258 0.5618 0.4309 0.5413 minicpm-llama2-v2.5 0.2651 0.3121 0.6246 0.6093 minigpt4_llama_2 0.1224 0.0738 0.5841 0.8019 minigpt4_vicuna-13b 0.0757 0.0460 0.2390 0.2701 minigpt_v2 0.5122 0.3530 0.5453 0.6974 mplug-owl2 0.1906 0.1406 0.8498 0.8805 otter 0.6371 0.2340 0.4989 0.8728 phi-3-vision 0.3599 0.1599 0.8534 0.9046 qwen-vl-chat 0.2882 0.1503 0.7734 0.8644 shikra-7b 0.1184 0.0540 0.4287 0.3008 yi-vl 0.0121 0.0231 0.9254 0.9254 GPT-4o 0.5987 0.6306 0.7971 0.7525 Gemini-l.5-pro 0.5101 0.6446 0.8584 0.7673
Table 9: Privacy Question Detection results before and after the addition of safety prompts. ACC is reported for evaluation. “w/o” stands for “without safety prompts” and “w/” stands for “with safety prompts”.
From Table 9, we find that the influence of safety prompts on models’ privacy awareness exhibits two distinct patterns: positive enhancement and inverse enhancement.
In the positive enhancement pattern, the changes in model test results mirror the performance in Privacy Leakage. Upon the introduction of safety prompts, some models demonstrate an increased recognition rate for sensitive questions but a decreased rate for insensitive ones. Notably, the change in privacy leakage is more pronounced than the change in privacy awareness after the incorporation of safety prompts. For instance, GPT-4o [33]’s $RtA$ for sensitive questions increased by 0.14 after adding safety prompts (see Table 10), while the recognition rate for sensitive questions only improved by 0.032.
In the inverse enhancement pattern, we observe that incorporating safety prompts reduces the recognition rate of sensitive questions for some models, like otter [20] and phi [1], while increasing their recognition rate for insensitive questions. However, this inverse enhancement does not negatively impact the models’ performance in Privacy Leakage. For example, although the addition of safety prompts decreased phi [1]’s recognition rate for sensitive questions by 0.20, its $RtA$ for sensitive questions actually improved by 0.18. The existence of this inverse enhancement pattern further underscores a significant disconnect between awareness and behavior in current models.
D.4 Other detailed results of Multi-$\textbf{P}^{\textbf{2}}$
Model Sensitive Questions Insensitive Questions w/o w/ w/o w/ blip2-opt-3b 0.1167 0.3452 0.9561 0.7957 blip2-opt-7b 0.0552 0.2819 0.9810 0.7774 blip2_flan-t5-xl 0.0223 0.0673 0.9946 0.9887 glm-4v-9b 0.5341 0.5568 0.7295 0.6776 instructblip_flan-t5-xl 0.0827 0.1004 0.9682 0.9715 instructblip_flan-t5-xxl 0.1119 0.0100 0.9455 0.9978 instructblip_vicuna-13b 0.1154 0.1459 0.9434 0.9376 instructblip_vicuna-7b 0.1792 0.1983 0.9463 0.9726 internlm-xcomposer-vl-7b 0.1972 0.2019 0.9309 0.9163 llava_1.5-13b 0.2362 0.4280 0.9228 0.8530 llava_1.5-7b 0.1856 0.3850 0.9495 0.8819 minicpm-llama2-v2.5 0.3867 0.6968 0.8147 0.6876 minigpt4_llama_2 0.3038 0.7421 0.9148 0.7292 minigpt4_vicuna-13b 0.3719 0.5716 0.8063 0.6864 minigpt_v2 0.1242 0.2289 0.9579 0.9470 mplug-owl2 0.2002 0.5003 0.9276 0.7054 otter 0.2068 0.1389 0.9266 0.9628 phi-3-vision 0.7017 0.8892 0.6535 0.4572 qwen-vl-chat 0.0485 0.2084 0.9736 0.7683 shikra-7b 0.1085 0.1446 0.9144 0.9043 yi-vl 0.2975 0.4496 0.8962 0.8268 GPT-4o 0.5414 0.6885 0.7415 0.6154 Gemini-1.5-pro 0.3797 0.5204 0.7651 0.6711
Table 10: Detailed results on Privacy Leakage before and after the addition of safety prompts. “w/o” stands for “without safety prompts” and “w/” stands for “with safety prompts”. The metrics for Sensitive and Insensitive Questions are RtA and 1 - RtA, respectively.
Model Sensitive Questions Insensitive Questions $ACC$ $RtA$ $ACC$ $1-RtA$ blip2-opt-3b 0.6486 0.1166 0.2507 0.9561 blip2-opt-7b 0.2650 0.0556 0.2039 0.9810 blip2_flan-t5-xl 0.0071 0.0223 0.9254 0.9945 glm-4v-9b 0.0446 0.5341 0.9203 0.7295 instructblip_flan-t5-xl 0.0116 0.0826 0.9254 0.9682 instructblip_flan-t5-xxl 0.0406 0.1118 0.9254 0.9455 instructblip_vicuna-13b 0.1055 0.1154 0.8849 0.9433 instructblip_vicuna-7b 0.4865 0.1791 0.5303 0.9463 internlm-xcomposer-vl-7b 0.0354 0.1972 0.9243 0.9309 llava_1.5-13b 0.4676 0.2361 0.6488 0.9229 llava_1.5-7b 0.6258 0.1856 0.4309 0.9496 minicpm-llama2-v2.5 0.2651 0.3868 0.6246 0.8147 minigpt4_llama_2 0.1224 0.3038 0.5841 0.9148 minigpt4_vicuna-13b 0.0757 0.3718 0.2390 0.8063 minigpt_v2 0.5122 0.1243 0.5453 0.9580 mplug-owl2 0.1906 0.2002 0.8498 0.9276 otter 0.6371 0.2068 0.4989 0.9265 phi-3-vision 0.3599 0.7017 0.8534 0.6535 qwen-vl-chat 0.2882 0.0485 0.7734 0.9737 shikra-7b 0.1184 0.1085 0.4287 0.9148 yi-vl 0.0121 0.2975 0.9254 0.8962 GPT-4o 0.5987 0.5414 0.7971 0.7415 Gemini-l.5-pro 0.5101 0.3797 0.8584 0.7651
Table 11: Detailed results on Sensitive Questions and Insensitive Questions. Sensitive Questions is the combination of samples in Perception Leakage, Reasoning Leakage, and Memory Leakage. $ACC$ reflects the results on Privacy Question Detection, while $RtA$ and $1-RtA$ represent the results on Privacy Leakage.
We evaluate 21 open-source models and 2 closed-source models on Multi-$\text{P}^{\text{2}}$. Table 11 and Table 10 corresponds the complete results of Section 5.5 and Section 5.6, respectively. Table 11 shows the detailed results on Privacy Leakage and Privacy Question Detection which reflect the mismatch between model’s privacy awareness and its privacy safeguard. Table 10 exhibits the impact of safety prompts on model’s privacy preservation capacities.
Appendix E Samples of test case
@@
In this section, we primarily showcase samples of Multi-$\text{P}^{\text{2}}$, covering as many testing perspectives as possible included in our benchmark. Fig 5 exhibits samples of three tasks in Privacy Awareness and Fig 6 shows samples in Privacy Leakage across three privacy types and three categories of privacy leakage.

### v3

@@ Appendix A Discussion
A-A Construction of Sensitive Questions
@@
A-B Copyright Attribution for Images
@@
TABLE VIII: Results of GPT-4o and human on Multi-P2A.
Eval Privacy Image Recognition Privacy Question Detection Perception Leakage Reasoning Leakage Memory Leakage Insensitive Questions ACCACC ACCACC RtARtA RtARtA RtARtA 1RtA1-RtA GPT-4o 0.750 0.698 0.476 0.478 0.671 0.742 human 0.943 0.958 0.964 0.979 0.951 0.936
A-C Quality analysis of the dataset
To evaluate data quality, we conduct user study involving 3 computer science experts, as shown in Table VIII. We randomly select 1,000 samples from Privacy Image Recognition and 1,000 samples from Privacy Question Detection, respectively. In the evaluation of Privacy Awareness, we directly provide each expert with samples from ours benchmark and require them to answer these closed-ended questions. Specifically, for Privacy Image Recognition and Privacy Question Detection, experts will evaluate the privacy sensitivity of samples by selecting Yes/No. Regarding Privacy Leakage, since all samples in this task have a one-to-one correspondence with those in Privacy Question Detection, and for human evaluators, considering a question to involve private information is equivalent to refusing to answer it, we directly used the human evaluation results from Privacy Question Detection as a substitute for assessing the Privacy Leakage. We find that humans achieve superior performance across tasks, while models like GPT-4o underperform. This demonstrates our data’s high accuracy and effectiveness in exposing model privacy risks.
A-D Concerns about Privacy Image Recognition
@@ Appendix B Information about Dataset and Models
B-A Dataset Statistics
In total, Multi-PA comprises 31,962 samples. Statistics for each task in Multi-PA are presented in Table IX.
TABLE IX: Statistics of Multi-$\text{P}^{\text{2}}$A.
@@ B-A Dataset Statistics
B-B Model Hub
Configurations and specifications for some open-source models evaluated in our experiments are listed in Table X.
TABLE X: Configurations of some open-source LVLMs evaluated in our experiments, “VE” stands for visual encoder and “LLM” stands for language model.
model VE LLM blip2-opt-3b [li2023blip] ViT-g/14(EVA-CLIP) OPT-3B blip2-opt-7b [li2023blip] ViT-g/14(EVA-CLIP) OPT-7B blip2_flan-t5-xl [li2023blip] ViT-g/14(EVA-CLIP) FlanT5-XL glm-4v-9b [glm2024chatglm] EVA-02-CLIP-E GLM-4-9B-Chat instructblip_flan-t5-xl [dai2023instructblipgeneralpurposevisionlanguagemodels] ViT-g/14(EVA-CLIP) FlanT5-XL instructblip_flan-t5-xxl [dai2023instructblipgeneralpurposevisionlanguagemodels] ViT-g/14(EVA-CLIP) FlanT5-XXL instructblip_vicuna-13b [dai2023instructblipgeneralpurposevisionlanguagemodels] ViT-g/14(EVA-CLIP) Vicuna-13B instructblip_vicuna-7b [dai2023instructblipgeneralpurposevisionlanguagemodels] ViT-g/14(EVA-CLIP) Vicuna-7B internlm-xcomposer-vl-7b [internlmxcomposer] ViT-g/14(EVA-CLIP) internlm-xcomposer-7b llava_1.5-13b [liu2024visual] ViT-L/14-336(CLIP) Vicuna-13B llava_1.5-7b [liu2024visual] ViT-L/14-336(CLIP) Vicuna-7B minicpm-llama2-v2.5 [yao2024minicpm] SigLIP SoViT-400M Llama3-Instruct-8B minigpt4_llama2 [zhu2023minigpt] BLIP2-VE(EVA-CLIP) LLaMA2 minigpt4_vicuna-13b [zhu2023minigpt] BLIP2-VE(EVA-CLIP) Vicuna-13B minigpt_v2 [chen2023minigpt] ViT-g/14(EVA-CLIP) LLaMA2 mplug-owl2 [ye2023mplugowl2] VIT-L(CLIP) LLaMA otter [li2023ottermultimodalmodelincontext] ViT-L/14(CLIP) LLaMA phi-3-vision [abdin2024phi] ViT-L/14-336(CLIP) phi-3 qwen-vl-chat [Qwen-VL] ViT-G(CLIP) Qwen-7b shikra-7b [chen2023shikra] ViT-L/14(CLIP) Vicuna-7B yi-vl[ai2024yi] OpenCLIP ViT-H yi-6B-chat
@@ Appendix C Details of dataset generation
C-A Detailed Privacy Categories
@@
C-B Questions of Privacy Question Detection
@@
C-C Questions of Privacy InfoFlow Assessment
@@
C-D Templates of Safety Enhancement
@@
C-E Pearson correlation in Privacy InfoFlow Assessment
Each sample receives two scores: model score $x$ and human score $y$. To get model score, models are required to select one score from five options:-100) strongly disagree, -50) somewhat disagree, 0) Neutral, 50) somewhat agree, 100) strongly agree. The human scores are inherited from CONFAIDE’s annotations. Pearson correlation coefficient is calculated based on model score set of all samples $X=\{x_{0},...,x_{391}\}$ and human score set $Y=\{y_{0},...,y_{391}\}$.
C-F Samples of test case
@@
In this section, we primarily showcase samples of Multi-PA, covering as many testing perspectives as possible included in our benchmark. Fig 6 exhibits samples of three tasks in Privacy Awareness and Fig 7 shows samples in Privacy Leakage across three privacy types and three categories of privacy leakage.